A 10 gram mixture of C2H6 and C3H4 gases is combusted.During the reaction, 30.8 gram of CO2 gas is produced.Find the mass of C2H6 in the initial mixture.
1
Expert's answer
2017-12-21T03:55:54-0500
Answer:
Let x = g C 3 H 4
and y = g C 2 H 6
equation 1 is x + y = 10.00 g
C 3 H 4 + 4O 2 ==> 3CO 2 + 2H 2 O
2C 2 H 6 + 7O 2 ==> 4CO 2 + 6H 2 O
Converting g C 3 H 4 to g CO 2 .
x(3*molar mass CO 2 /molar mass C 3 H 4 )
Converting g C 2 H 6 to g CO 2
y(4*molar mass CO 2 /2*molar mass C 2 H 6 )
The total is 30.8 g so put all of that together for equation 2 of
x(3*molar mass CO 2 /molar mass C 3 H 4 ) + y(4*molar mass CO 2 /2*molar mass C 2 H 6 ) -= 30.8
M (CO 2 ) = 44 g/mol
M (C 3 H 4 ) = 40 g/mol
M (C 2 H 6 ) = 30 g/mol
x(3*molar mass CO 2 /molar mass C 3 H 4 ) + y(4*molar mass CO 2 /2*molar mass C 2 H 6 ) -= 30.8
3.3 x + 2.9 y = 30.8
x + y = 10.00
x = 10-y
3.3 (10-y) + 2.9 y = 30.8
33-3.3y + 2.9 y = 30.8
33-0.4y = 30.8
-0.4 y = -2.2
Y = 5.5 g
m (C 2 H 6 ) = 5.5 g
m (C 3 H 4 ) = 10-5.5 = 4.5 g
Now solving the two equations simultaneously and solving for x = grams C 3 H 4 .